OpenStudy (anonymous):

Precalculus: What is discontinuity in rational functions? Is it another way of saying there is a hole in the graph? Discontinuity: Anytime the denominator of a rational function equals zero. What if it's 0/0? Is it a discontinuity or a hole? Or is it both? Does it have to be a number over 0 to count as a discontinuity?

OpenStudy (anonymous):

No, it is still a discontinuity. Just imagine the function f(x) = x/x does the point x=0 exist? Nope.

OpenStudy (anonymous):

so 0/0 is not a hole?

OpenStudy (anonymous):

It is a hole. The numerator does no have to be over 0 to count as a discontinuity.

OpenStudy (anonymous):

How does it count as a discontinuity?

OpenStudy (anonymous):

You cannot divide by zero, therefore the point does not exist

OpenStudy (anonymous):

it is a jump discontinuity

OpenStudy (anonymous):

Wait, isn't that what a vertical asymptote is?

OpenStudy (anonymous):

Do you have a graphing calculator?

OpenStudy (anonymous):

or just a regular calculator?

OpenStudy (anonymous):

graphing calculator

OpenStudy (anonymous):

Graph the function y = x/x then go to >2nd>calc>value x=0

OpenStudy (anonymous):

It will give you an error because the point does not exist, it is 0/0. And it is a jump discontinuity. The condition for the discontinuity is that the denominator is 0, the numerator DOES NOT matter

OpenStudy (anonymous):

You are confusing discontinuities with limits, where 0/0 could have an actual value (limit)

ganeshie8 (ganeshie8):

i guess you wanto knw how to knw whether a discontinuity has vertical asymptote or not

OpenStudy (anonymous):

Wait, so in f(x) = 1/(3x^2+3x-18) The vertical asymptotes are 2,-3 which are also the discontinuities?

OpenStudy (anonymous):

x=2 and x=-3

OpenStudy (anonymous):

are vertical asymptotes and discontinuities the same thing?

OpenStudy (anonymous):

vertical asymptotes are discontinuities but discontinuities are not necessarily vertical asymptotes

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

Why can't discontiunities be vertical aympsotes?

ganeshie8 (ganeshie8):

sometimes the factors leading to 1/0 cancel out in numerator and denum (just the case with y=x/x) leaving a discontinuity(hole) without any asymptote, cuz around the hole, graph stays at fixed value 1 here, it doesnt increase forever...

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

consider another example where factors leading to 1/0 doesnt cancel out. y = 1/x

ganeshie8 (ganeshie8):